OpenStudy (anonymous):

Solve the equation using the Zero Product Property. (2x – 4)(2x – 1) = 0

OpenStudy (anonymous):

2x-4=0 or 2x+1=0 2x=4 or 2x=-1 x=2 or x=-1/2

OpenStudy (anonymous):

thank you!